# Classical harmonic analysis viewed through the prism of noncommutative geometry

### v3

@@ 8 Examples of triples and Banach Fredholm modules
Suppose that $2\leqslant p<\infty$. Consider the triple $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{p}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$. The operator $\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is bisectorial of angle 0 and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\omega}^{\mathrm{bi}})$ functional calculus for any angle $\omega>0$ and the commutator $\big[\tfrac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x},f\big]$ is a well-defined bounded operator for any function $f\in\mathrm{C}^{\infty}_{0}(\mathbb{R})$.
@@
Proof : Note that it is well-known [EnN00, Proposition 1 p. 66] that the operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ generate a strongly continuous group of operators acting on $\mathrm{L}^{p}(\mathbb{R})$, namely the group of translations. By Example 2.7 and Example 2.8, we deduce that the unbounded operator $\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}\theta}$ is bisectorial of angle 0 and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\omega}^{\mathrm{bi}})$ functional calculus for any angle $\omega>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. For any function $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$, an elementary computation reveals that $M_{f}\cdot\operatorname{dom}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}\subset\operatorname{dom}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ and that
@@
(8.16) $\big[\tfrac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x},f\big]=\tfrac{1}{\mathrm{i}}\big(\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}M_{f}-M_{f}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}\big)=\frac{1}{\mathrm{i}}M_{f^{\prime}}.$
So $\frac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x},f]$ defines a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{R})$.
@@
and can be seen by [HvNVW16, Proposition 5.2.2 p. 389] as a Fourier multiplier with symbol $\mathrm{i}\,\operatorname{\mathrm{sign}}(\xi)$, i.e. we have
@@
Note that it was observed in [Con94, p. 314] that $\mathrm{i}\mathcal{H}$ is a Fredholm module on the Hilbert space $\mathrm{L}^{2}(\mathbb{R})$. Note that the commutator is defined by (1.6)
For the computation of the index pairing, we need some information on Hardy spaces. Recall the notation $\mathbb{C}_{+}\overset{\mathrm{def}}{=}\{z\in\mathbb{C}:\mathrm{Im}z>0\}$. Following [Nik02, Definition 6.3.2 p. 145], we denote by $\mathrm{H}^{p}(\mathbb{C}_{+})$ the space of all functions $F$ which are analytic in the upper half plane $\mathbb{C}_{+}$ such that
@@
of the Banach space $\mathrm{L}^{p}(\mathbb{R})$. Cauchy’s representation theorem in Hardy spaces, e.g. [Mas09, Chapter 13], can be formulated as follows. Consider some analytic function $F\colon\mathbb{C}_{+}\to\mathbb{C}$. Then the following assertions are equivalent.
@@
    The function $F$ belongs to the space $\mathrm{H}^{p}(\mathbb{C}_{+})$.
@@
    There exists a unique function $f\in\mathrm{H}^{p}(\mathbb{R})$ such that
@@
In this case, we have
$\left\|F\right\|_{\mathrm{H}^{p}(\mathbb{C}_{+})}=\left\|f\right\|_{\mathrm{L}^{p}(\mathbb{R})}.$
and the non-tangential boundary function of $F$ is equal to $f$, i.e. $f(x)=\lim_{z\to x,z\in\mathbb{C}_{+}}F(z)$. Consequently, we have an isometric isomorphism of the Hardy space $\mathrm{H}^{p}(\mathbb{C}_{+})$ onto the space $\mathrm{H}^{p}(\mathbb{R})$. Thus the Hardy space $\mathrm{H}^{p}(\mathbb{C}_{+})$ can be identified with a closed subspace of the Banach space $\mathrm{L}^{p}(\mathbb{R})$.
The Toeplitz operator $T_{g}\colon\mathrm{H}^{p}(\mathbb{R})\to\mathrm{H}^{p}(\mathbb{R})$ with symbol $g\in\mathrm{L}^{\infty}(\mathbb{R})$ is defined by
(8.20) $T_{g}(f)\overset{\mathrm{def}}{=}P_{+}(gf),\quad f\in\mathrm{H}^{p}(\mathbb{R}).$
If $f\colon\mathbb{R}\cup\{\infty\}\to\mathbb{C}$ is a continuous function then the image of $f$ in the complex plane is a closed curve $\Gamma$. If
(8.21) $\inf\limits_{x\in\mathbb{R}\cup\{\infty\}}|f(x)|>0,$
we define the winding number $\operatorname{wind}f$ as the winding number of $\Gamma$ around the origin. The condition (8.21) is equivalent to say that $f$ is invertible in the algebra $\mathrm{C}(\mathbb{R}\cup\{\infty\})$.
We will use the following result [Cam17, Theorem 6.2].
@@
Let $f\in\mathrm{C}(\mathbb{R}\cup\{\infty\})$. The operator $T_{f}\colon\mathrm{H}^{p}(\mathbb{R})\to\mathrm{H}^{p}(\mathbb{R})$ has closed range if and only if (8.21) is satisfied. In this case, the operator $T_{f}$ is Fredholm and its Fredholm index is $\operatorname{Index}T_{f}=-\operatorname{wind}f$.
Remark 8.11
It is known that the condition (8.21) is equivalent to the existence of a Wiener-Hopf $p$-factorization of $f$. In this case, $f=f_{-}r^{k}f_{+}$ with $k=\operatorname{wind}f$.
The following observation is new even if $p=2$. Recall that $\mathrm{K}_{1}(\mathrm{C}_{0}(\mathbb{R}))=\mathbb{Z}$, see e.g. [WeO93].

### v4

@@ 8 Examples of triples and Banach Fredholm modules
Let $1<p<\infty$ and let $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathrm{d}}{\mathrm{d}x}$ with domain $\operatorname{dom}D=\mathrm{W}^{1,p}(\mathbb{R})$. Then
(8.20) $\mathrm{W}^{1,\infty}(\mathbb{R})=\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{R})).$
Moreover, for any function $f\in\mathrm{W}^{1,\infty}(\mathbb{R})$ we have
(8.21) $D,M_{f}]=\frac{1}{\mathrm{i}}M_{f^{\prime}}\quad\text{and}\quad\left\|[D,M_{f}]\right\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}=\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}.$
Proof : If $f\in\mathrm{W}^{1,\infty}(\mathbb{R})$ then $M_{f}(\mathrm{W}^{1,p}(\mathbb{R}))\subset\mathrm{W}^{1,p}(\mathbb{R})$ by [BeH21, Theorem 4.1 p. 284]. Moreover, if $u\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ we have
$D,M_{f}]u=(DM_{f}-M_{f})u=\tfrac{1}{\mathrm{i}}\big((fu)^{\prime}-fu^{\prime}\big)=\tfrac{1}{\mathrm{i}}f^{\prime}u.$
Hence $D,M_{f}]=\tfrac{1}{\mathrm{i}}M_{f^{\prime}}$ extends boundedly on $\mathrm{L}^{p}(\mathbb{R})$ with $\left\|[D,M_{f}]\right\|_{\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})}=\left\|f^{\prime}\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}$.
Let $(T_{t})_{t\in\mathbb{R}}$ be the translation group on $\mathrm{L}^{p}(\mathbb{R})$ given by $(T_{t}u)(x)=u(x+t)$. Then for any $t\in\mathbb{R}$ we have $T_{t}=\mathrm{e}^{\mathrm{i}tD}$ (see [EnN00, Proposition 1 p. 66] or [Haa06, Proposition 8.4.2 p. 229]) and
(8.22) $T_{t}M_{f}T_{-t}=M_{f(\cdot+t)},\qquad t\in\mathbb{R}.$
Let $f\in\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{R}))$. Since $D,M_{f}]\in\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))$ by Proposition 8.2 we deduce that
[…]
If $|t|$ is small, we infer that
[…]
$\leqslant\int_{0}^{1}\left\|T_{st}(\mathrm{i}[D,M_{f}])T_{-st}\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}\mathop{}\mathopen{}\mathrm{d}s\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}.$
By Banach–Alaoglu, there exist $t_{\alpha}\to 0$ and $g\in\mathrm{L}^{\infty}(\mathbb{R})$ such that $\Delta_{t_{\alpha}}f\overset{}{\longrightarrow}g$ in the weak* topology of the dual space $\mathrm{L}^{\infty}(\mathbb{R})$, with $\left\|g\right\|_{\mathrm{L}^{\infty}(\mathbb{R})}\leqslant\left\|[D,M_{f}]\right\|_{\mathrm{B}(\mathrm{L}^{p}(\mathbb{R}))}$. For any function $\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ we have
$\int_{\mathbb{R}}(\Delta_{t_{\alpha}}f)\varphi=\int_{\mathbb{R}}f\frac{\varphi(\cdot-t_{\alpha})-\varphi}{t_{\alpha}}\longrightarrow-\int_{\mathbb{R}}f\,\varphi^{\prime}$
since $\frac{\varphi(\cdot-t)-\varphi}{t}\to-\varphi^{\prime}$ in $\mathrm{L}^{1}(\mathbb{R})$ as $t\to 0$. Passing to the weak* limit gives
$\int_{\mathbb{R}}g\varphi=-\int_{\mathbb{R}}f\varphi^{\prime},\quad\varphi\in\mathrm{C}_{c}^{\infty}(\mathbb{R}),$
i.e., $g$ is the weak derivative of $f$. Hence $f\in\mathrm{W}^{1,\infty}(\mathbb{R})$ and $f^{\prime}=g\in\mathrm{L}^{\infty}(\mathbb{R})$.
Proposition 8.13
@@
Suppose $1<p<\infty$. Then the operator $D=\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ on $\mathrm{L}^{p}(\mathbb{R})$ with domain $\mathrm{W}^{1,p}(\mathbb{R})$ is bisectorial of angle 0, i.e., $\omega_{\mathrm{bi}}(D)=0$, and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus for any angle $\theta>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$.
@@
Proof : The operator $\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ generates a strongly continuous group of operators acting on the Banach space $\mathrm{L}^{p}(\mathbb{R})$, namely the left translation group, defined by $(T_{t}g)(x)=g(x+t)$ where $t,x\in\mathbb{R}$ (see [EnN00, Proposition 1 p. 66] or [Haa06, Proposition 8.4.2 p. 229]). By Example 2.7 and Example 2.8, we deduce that the unbounded operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ is bisectorial with $\omega_{\mathrm{bi}}(D)=0$ and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus for any angle $\theta>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$.
Proposition 8.14
Suppose $1<p<\infty$. Consider the operator $D=\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ on $\mathrm{L}^{p}(\mathbb{R})$ with domain $\mathrm{W}^{1,p}(\mathbb{R})$. For any function $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ the operator $M_{f}(\mathrm{i}+D)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$ is compact.
Proof : First proof. Consider the infinitesimal generator $A\overset{\mathrm{def}}{=}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ of the left translation group $(T_{t})_{t\in\mathbb{R}}$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$. By the Laplace formula [HvNVW18, Proposition G.4.1 p. 532] for the resolvent, we have for $\operatorname{Re}\lambda>0$ the equality
(8.24) $(\lambda-A)^{-1}g=\int_{0}^{\infty}\mathrm{e}^{-\lambda t}T_{t}g\mathop{}\mathopen{}\mathrm{d}t,\quad g\in\mathrm{L}^{p}(\mathbb{R}).$
Taking $\lambda=1$ and using $\mathrm{i}+D=\mathrm{i}(1-A)$, we obtain for almost all $x\in\mathbb{R}$
[…]
$\overset{t=y-x}{=}-\mathrm{i}\int_{x}^{\infty}\mathrm{e}^{x-y}g(y)\mathop{}\mathopen{}\mathrm{d}y=-\mathrm{i}\int_{\mathbb{R}}k(x-y)g(y)\mathop{}\mathopen{}\mathrm{d}y,$
with $k(t)\overset{\mathrm{def}}{=}\mathrm{e}^{t}1_{(-\infty,0]}(t)$. Hence $M_{f}(\mathrm{i}+D)^{-1}$ is an integral operator with kernel
$K(x,y)\overset{\mathrm{def}}{=}-\mathrm{i}f(x)k(x-y).$
Note that $k\in\mathrm{L}^{p^{*}}(\mathbb{R})$, where $p^{\ast}$ is the conjugate exponent defined by $\frac{1}{p}+\frac{1}{p^{*}}=1$. For almost all $x\in\mathbb{R}$, we have
(8.26) $\left\|y\mapsto K(x,y)\right\|_{\mathrm{L}^{p^{*}}(\mathbb{R})}=|f(x)|\left\|y\mapsto k(x-y)\right\|_{\mathrm{L}^{p^{*}}(\mathbb{R})}=|f(x)|\left\|k\right\|_{\mathrm{L}^{p^{*}}(\mathbb{R})},$
Consequently, we have
[…]
since $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$. The condition $K\in\mathrm{L}^{p}(\mathbb{R},\mathrm{L}^{p^{*}}(\mathbb{R}))$ implies that $M_{f}(\mathrm{i}+D)^{-1}$ is a Hille-Tamarkin operator, hence compact on $\mathrm{L}^{p}(\mathbb{R})$ by [Jor82, Theorem p. 446] or [HiT34, Theorem 11.6 p. 275].
Second proof. The resolvent operator $(\mathrm{i}+D)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$ is bounded by definition. Let $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$. Consider a bounded open subset $\Omega\subset\mathbb{R}$ such that $\operatorname{supp}f\subset\Omega$. Note that we have a bounded operator $M_{f}\colon\mathrm{W}^{1,p}(\mathbb{R})\to\mathrm{W}^{1,p}(\Omega)$. We denote by $J\colon\mathrm{W}^{1,p}(\Omega)\to\mathrm{L}^{p}(\Omega)$ the canonical injection, which is compact by [Eva10, Remark p. 289]. If $\mathcal{E}\colon\mathrm{L}^{p}(\Omega)\to\mathrm{L}^{p}(\mathbb{R})$ is the extension map by 0, then the operator $M_{f}(\mathrm{i}+D)^{-1}$ admits the factorization
$M_{f}(\mathrm{i}+D)^{-1}=E\circ J\circ M_{f}\circ(\mathrm{i}+D)^{-1},$
Since this factorization contains a compact operator, it is also a compact operator.
Actually, we have the following quantitative compactness result relying on the Poincaré inequality [EdE23, Theorem 2.4 p. 21]
(8.27) $\left\|u-\frac{1}{|\Omega|}\int_{\Omega}u\right\|_{\mathrm{L}^{p}(\Omega)}\leqslant\bigg(\frac{\omega_{d}}{|\Omega|}\bigg)^{1-\frac{1}{d}}\operatorname{diam}(\Omega)^{d}\left\|\nabla u\right\|_{\mathrm{L}^{p}(\Omega,\ell^{2}_{d})}$
for a bounded convex subset $\Omega$ of $\mathbb{R}^{d}$.
Proposition 8.15
Suppose that $1<p<\infty$. For any function $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ the operator $M_{f}(D+\mathrm{i})^{-1}$ belongs to $S_{\mathrm{app}}^{1,\infty}(\mathrm{L}^{p}(\mathbb{R}))$.
Proof : Since $D^{2}=-\Delta$, we have the algebraic identity
$(D+\mathrm{i})^{-1}=-(\mathrm{i}-D)(1-\Delta)^{-1}.$
Note that the Bessel potential operator $(1-\Delta)^{-1}$ induces a bounded operator $(1-\Delta)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{2,p}(\mathbb{R})$, while the operator $\mathrm{i}-D\colon\mathrm{W}^{2,p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$ is bounded by [Kry08, Exercise 8 p. 10]. Hence the operator
$S\overset{\mathrm{def}}{=}(\mathrm{i}-D)(1+\Delta)^{-1}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{W}^{1,p}(\mathbb{R})$
is bounded. Let $I$ be a fixed bounded interval of $\mathbb{R}$ containing $\operatorname{supp}f$. We consider the restriction operator $R_{I}\colon\mathrm{W}^{1,p}(\mathbb{R})\to\mathrm{W}^{1,p}(I)$ and the extension by zero operator $E_{I}\colon\mathrm{L}^{p}(I)\to\mathrm{L}^{p}(\mathbb{R})$. If $J\colon\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)$ is the Sobolev embedding on the bounded interval $I$, then we have the factorization
(8.28) $M_{f}(D+\mathrm{i})^{-1}=E_{I}\underbrace{M_{f}|_{\mathrm{L}^{p}(I)}}_{M_{f}^{I}\overset{\mathrm{def}}{=}}\underbrace{J}_{\mathrm{W}^{1,p}(I)\hookrightarrow\mathrm{L}^{p}(I)}\underbrace{R_{I}S}_{S_{I}\overset{\mathrm{def}}{=}}.$
Consider a partition of $I$ into $n$ subintervals $I_{1},\ldots,I_{n}$ of equal length $\frac{|I|}{n}$. Consider the linear operator $P_{n}\colon\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)$ defined by
$(P_{n}g)(x)\overset{\mathrm{def}}{=}\sum_{k=1}^{n}\bigg(\frac{1}{|I_{k}|}\int_{I_{k}}g(t)\mathop{}\mathopen{}\mathrm{d}t\bigg)1_{I_{k}}(x),\qquad g\in\mathrm{W}^{1,p}(I).$
Observe that $\operatorname{rank}P_{n}\leqslant n$. On each subinterval $I_{k}$, the Poincaré inequality yields
[…]
Summing over $k=1,\dots,n$ (the $I_{k}$ are disjoint) gives
@@
$\left\|g-P_{n}(g)\right\|_{\mathrm{L}^{p}(I)}^{p}=\sum_{k=1}^{n}\left\|g-\frac{1}{|I_{k}|}\int_{I_{k}}g(t)\mathop{}\mathopen{}\mathrm{d}t\right\|_{\mathrm{L}^{p}(I_{k})}^{p}\leqslant\bigg(\frac{|I|}{n}\bigg)^{p}\sum_{k=1}^{n}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I_{k})}^{p}=\bigg(\frac{|I|}{n}\bigg)^{p}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I)}^{p}.$
Taking the $p$-th root, we obtain
$\left\|g-P_{n}g\right\|_{\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}\left\|g^{\prime}\right\|_{\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}\left\|g\right\|_{\mathrm{W}^{1,p}(I)}.$
Hence, we obtain the estimate
$\left\|J-P_{n}\right\|_{\mathrm{W}^{1,p}(I)\to\mathrm{L}^{p}(I)}\leqslant C\frac{|I|}{n}.$
By the ideal property of approximation numbers, this implies
[…]
From the previous factorization and the ideal property of approximation numbers,
[…]
We conclude with (5.9).
Theorem 8.16
Suppose that $1<p<\infty$. Then $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{p}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$ is a locally compact spectral triple which is $1^{+}$-summable, hence $q$-summable for any $q>1$. More precisely, we have
$\dim\big(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{p}(\mathbb{R}),\tfrac{1}{\mathrm{i}}\tfrac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}\big)=1.$
Proof : By Proposition 8.13, the unbounded operator $D\overset{\mathrm{def}}{=}\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x}$ is bisectorial and admits a bounded $\mathrm{H}^{\infty}(\Sigma_{\theta}^{\mathrm{bi}})$ functional calculus for any angle $\theta>0$ on the Banach space $\mathrm{L}^{p}(\mathbb{R})$.
By Proposition 8.12, the Lipschitz algebra $\operatorname{\mathrm{Lip}}_{D}(\mathrm{C}_{0}(\mathbb{R}))=\operatorname{\mathrm{Lip}}_{D}(\mathrm{L}^{\infty}(\mathbb{R}))\cap\mathrm{C}_{0}(\mathbb{R})$ contains $\mathrm{C}_{c}^{\infty}(\mathbb{R})$. Hence $\operatorname{\mathrm{Lip}}_{D}(\mathrm{C}_{0}(\mathbb{R}))$ is dense in $\mathrm{C}_{0}(\mathbb{R})$. Finally, Proposition 8.14 gives the compactness condition. The spectral triple is $1^{+}$-summable by Proposition 8.15.
Let $f\in\mathrm{C}_{c}^{\infty}(\mathbb{R})$ such that $f=1$ on a closed interval $I=[a,b]$. Consider the restriction operator $R_{I}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}([a,b])$ and the extension by zero operator $E_{I}\colon\mathrm{L}^{p}([a,b])\to\mathrm{L}^{p}(\mathbb{R})$. For almost all $x\in\mathbb{R}$ and any $g\in\mathrm{L}^{p}(\mathbb{R})$, we have seen that
[…]
Hence for almost all $x\in[a,b]$ and $g\in\mathrm{L}^{p}([a,b])$,
$\big(R_{I}M_{f}(\mathrm{i}+D)^{-1}E_{I}g\big)(x)=-\mathrm{i}\int_{x}^{b}e^{x-y}g(y)\mathop{}\mathopen{}\mathrm{d}y.$
Consider the Hardy operator $H\colon\mathrm{L}^{p}([a,b])\to\mathrm{L}^{p}([a,b])$ defined by $(Vh)(x)\overset{\mathrm{def}}{=}\int_{x}^{b}h(y)\mathop{}\mathopen{}\mathrm{d}y$ and the multiplication operators $U\colon\mathrm{L}^{p}([a,b])\to\mathrm{L}^{p}([a,b])$, $g\mapsto(y\mapsto\mathrm{e}^{-y}g(y))$, and $W\colon\mathrm{L}^{p}([a,b])\to\mathrm{L}^{p}([a,b])$, $x\mapsto(x\mapsto\mathrm{e}^{x}h(x))$. Then
(8.30) $R_{I}M_{f}(\mathrm{i}+D)^{-1}E_{I}=-\mathrm{i}WHU.$
Since $U$ and $W$ are bounded isomorphisms on the Banach space $\mathrm{L}^{p}([a,b])$, the ideal property of approximation numbers yields
[…]
Thus it suffices to use the estimate $a_{n}(H)\gtrsim\frac{1}{n}$ of [EdL07, Theorem 3.3 p. 25]. By (5.9), we conclude that $(\mathrm{C}_{0}(\mathbb{R}),\mathrm{L}^{p}(\mathbb{R}),\frac{1}{\mathrm{i}}\frac{\mathop{}\mathopen{}\mathrm{d}}{\mathop{}\mathopen{}\mathrm{d}x})$ is not 1-summable.
@@
and can be seen by [HvNVW16, Proposition 5.2.2 p. 389] as a Fourier multiplier with symbol $\mathrm{i}\,\operatorname{\mathrm{sign}}(\xi)$, i.e., we have
@@
Note that it was observed in [Con94, p. 314] (see also [GVF01, p. 330] and [LMSZ23, p. 349]), that $F\overset{\mathrm{def}}{=}\mathrm{i}\mathcal{H}$ is a Fredholm module on the Hilbert space $\mathrm{L}^{2}(\mathbb{R})$. As observed in [GaR85, p. 473], the commutator is defined by the formula (1.6).
Remark 8.17
A famous result of Coifman, Rochberg and Weiss [CRW76], described in [GaR85, p. 473] and [Wic20, Theorem p. 2], says that the formula (1.6) induces a bounded operator on the Banach space $\mathrm{L}^{p}(\mathbb{R})$ if and only if the function $f$ belongs to the space $\mathrm{BMO}(\mathbb{R})$.
Remark 8.18
If $0<q<\infty$, a classical result of Peller (and Semmes for the case $q<1$), described in [Con94, p. 314], says that the commutator $F,M_{f}]$ is in the Schatten class $S^{q}(\mathrm{L}^{2}(\mathbb{R}))$ if and only if the function $f$ belongs to the Besov space $\mathrm{B}^{q}_{q,q}(\mathbb{R})$.
For the computation of the index pairing, we need some information on Hardy spaces. Recall the notation $\mathbb{C}_{+}\overset{\mathrm{def}}{=}\{z\in\mathbb{C}:\mathrm{Im}z>0\}$ for the open upper half plane. Following [Nik02, Definition 6.3.2 p. 145] and [Mas09, p. 247], we denote by $\mathrm{H}^{p}(\mathbb{C}_{+})$ the space of all functions $F$ which are holomorphic in the upper half plane $\mathbb{C}_{+}$ such that
@@
of the Banach space $\mathrm{L}^{p}(\mathbb{R})$, considered in [Mas09, (13.4) p. 282]. Cauchy’s representation theorem in Hardy spaces, e.g., [Mas09, Chapter 13], can be formulated as follows. For any holomorphic function $F\colon\mathbb{C}_{+}\to\mathbb{C}$, the following assertions are equivalent:
@@
    the function $F$ belongs to the space $\mathrm{H}^{p}(\mathbb{C}_{+})$,
@@
    there exists a function $f\in\mathrm{H}^{p}(\mathbb{R})$ such that
@@
In this case, $f$ is unique, we have $\left\|F\right\|_{\mathrm{H}^{p}(\mathbb{C}_{+})}=\left\|f\right\|_{\mathrm{L}^{p}(\mathbb{R})}$ and the non-tangential boundary function of $F$ is equal to $f$, i.e., $f(x)=\lim_{z\to x,z\in\mathbb{C}_{+}}F(z)$. Consequently, we have an isometric isomorphism of the Hardy space $\mathrm{H}^{p}(\mathbb{C}_{+})$ onto the space $\mathrm{H}^{p}(\mathbb{R})$. Thus the Hardy space $\mathrm{H}^{p}(\mathbb{C}_{+})$ can be identified with a closed subspace of the Banach space $\mathrm{L}^{p}(\mathbb{R})$. Recall that $P\overset{\mathrm{def}}{=}\frac{\mathrm{Id}+F}{2}\colon\mathrm{L}^{p}(\mathbb{R})\to\mathrm{L}^{p}(\mathbb{R})$ is a bounded projection on the subspace $\mathrm{H}^{p}(\mathbb{R})$.
Let $f\colon\mathbb{R}\cup\{\infty\}\to\mathbb{C}$ be a continuous function satisfying
(8.34) $\inf_{x\in\mathbb{R}\cup\{\infty\}}|f(x)|>0.$
Using the identification between one-point compactification $\mathbb{R}\cup\{\infty\}$ of $\mathbb{R}$ with the unit circle $\mathbb{T}$ via the standard homeomorphism $\Phi\colon\mathbb{R}\cup\{\infty\}\to\mathbb{T}$, $x\mapsto\frac{x-\mathrm{i}}{x+\mathrm{i}}$, we then define the winding number of $f$ by $\operatorname{wind}f\,\overset{\mathrm{def}}{=}\,\operatorname{wind}(f\circ\Phi^{-1})$. Note that condition (8.34) means that $f$ is invertible in the commutative Banach algebra $\mathrm{C}(\mathbb{R}\cup\{\infty\})$.
We will use the following result [Cam17, Theorem 6.2 p. 141].
@@
Let $f\in\mathrm{C}(\mathbb{R}\cup\{\infty\})$. The Toeplitz operator $T_{f}\colon\mathrm{H}^{p}(\mathbb{R})\to\mathrm{H}^{p}(\mathbb{R})$, $g\mapsto P(fg)$ with symbol $f$ has closed range if and only if (8.34) is satisfied. In this case, the operator $T_{f}$ is Fredholm and its Fredholm index is $\operatorname{Index}T_{f}=-\operatorname{wind}f$.
The following observation is new even if $p=2$. Recall that $\mathrm{K}_{1}(\mathrm{C}_{0}(\mathbb{R}))=\mathbb{Z}$, see for instance [WeO93, p. 123].